OpenStudy (anonymous):

HELP! PLEASE!!! Gary owns an island monkey preserve which has a length of 2x + 1 units and a width of x – 4 units. If the island is 1,586 square units in size, what are the dimensions of the island?

OpenStudy (anonymous):

(2x + 1)(x - 4) = 1586 Expand left hand side with FOIL. 2x^2 + x - 8x -4 = 1586 2x^2 - 7x -4 = 1586. Subtract 1586 from both sides: 2x^2 - 7x - 1590 = 0 Apply the quadratic formula (or use wolfram alpha). We get x = -53/2 or x = 30. x CANNOT be negative, so x = 30. 61 by 26 is your answer.